# When LLM-based Code Generation Meets the Software Development Process

### v1

@@ Abstract
Software process models play a pivotal role in fostering collaboration and communication within software teams, enabling them to tackle intricate development tasks effectively. This paper introduces LCG, a code generation framework inspired by established software engineering practices. LCG leverages multiple Large Language Model (LLM) agents to emulate various software process models, namely , , and . Each model assigns LLM agents specific roles such as requirement engineer, architect, developer, tester, and scrum master, mirroring typical development activities and communication patterns. Through collaborative efforts utilizing chain-of-thought and prompt composition techniques, the agents continuously refine themselves to enhance code quality. Utilizing GPT3.5 as the underlying LLM and baseline (GPT), we evaluate LCG across four code generation benchmarks: HumanEval, HumanEval-ET, MBPP, and MBPP-ET. Results indicate outperforms other models, achieving Pass@1 scores of 75.2, 65.5, 82.5, and 56.7 in HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively—an average 15% improvement over GPT. Analysis reveals distinct impacts of development activities on generated code, with design and code reviews contributing to enhanced exception-handling, while design, testing, and code reviews mitigate code smells. Furthermore, temperature values exhibit negligible influence on Pass@1 across all models. However, variations in Pass@1 are notable for different GPT3.5 model versions, ranging from 5 to over 60 in HumanEval, highlighting the stability of LCG across model versions. This stability underscores the importance of adopting software process models to bolster the quality and consistency of LLM-generated code.
@@
Given the complexity of software development, LLM agents stand out among various LLM techniques. Agents are LLM instances that can be customized to carry out specific tasks that replicate human workflow [12, 9]. Recently, multi-agent systems have achieved significant progress in solving complex problems in software development by emulating development roles [12, 9, 29]. MetaGPT, introduced by Hong et al. [12], integrated development workflow using standard operating procedures by assigning specific roles (e.g., a designer or a developer) to LLM agents. Dong et al. [9] developed self-collaboration, which assign LLM agents to work as distinct “experts” for sub-tasks in software development. Qian et al. [29] proposed an end-to-end framework for software development through self-communication among the agents.
@@
Despite the promising applications of LLMs in automating software engineering tasks, it is pivotal to recognize that software development is a collaborative and multi-faceted endeavor. In the real world, developers and stakeholders work together, following certain software process models like Waterfall, Test-Driven-Development (TDD), and Scrum. The process models help facilitate communication and collaboration to ensure the delivery of high-quality products. Even though there is a common community agreement on the pros and cons of each process model [11], the impact of adopting these process models for LLM code generation tasks remains unknown. In particular, will emulating different process models impact the generated code quality in different aspects, such as reliability, code smell, and functional correctness?
@@
While some research has explored integrating multi-agents within LLM frameworks [29, 43, 45], their research focus diverges from the influence of the software process model on code generations for several reasons: 1) Xu et al. [43] do not adhere to specific process models, and 2) both Dong et al. [9] and Qian et al. [29] focus solely on Waterfall, neglecting Test-Driven-Development (TDD) and Scrum, which may have different impact on code generations. Importantly, none of the aforementioned studies conduct a fine-grain analysis of how different development activities affect code quality metrics, such as code smell and reliability, other than the Pass@1 score. Our study takes further steps to analyze the impacts of different agents within the process models on code generation and their influence on other code quality attributes.
@@
This paper presents a novel multi-agent Large Language Model (LLM) based code generation framework named LCG. LCG integrates diverse prompt engineering techniques, including chain-of-thought [38, 18], prompt composition [20, 48], and self-refinement [23], with a focus on emulating the flow of development activities in various software process models. Specifically, we implemented three popular process models into LCG: , , and . Each process model emulates a real-world development team involving several LLM agents whose roles (i.e., requirement engineer, architect, developer, tester, and scrum master) correspond to common software development activities. The agents are assigned roles with different domain-specific expertise, working collaboratively to produce software artifacts and help other agents review and improve the artifacts in every activity of a process model.
@@
We evaluate LCG on four popular code generation benchmarks: HumanEval [6], HumanEval-ET [8], MBPP [2], and MBPP-ET [19]. We apply zero-shot learning to avoid biases in selecting few-shot samples [42]. To compare, we also apply zero-shot learning on GPT-3.5 as our baseline (GPT). We repeat our experiments five times to account for variability in LLM’s responses and report the average value and standard deviation. To study code quality, in addition to Pass@1, we run static code checkers to detect the prevalence of code smells in the generated code. Our evaluation shows that ’s generated code achieves the highest accuracy (Pass@1 is 75.2 for HumanEval, 65.5 for HumanEval-ET, 82.5 for MBPP, and 56.7 for MBPP-ET), surpassing GPT’s Pass@1 by 5.2% to 31.5%. While LCG, in general, is more stable than GPT, exhibits the most stable results with an average standard deviation of only 1.3% across all benchmarks.
@@
We further study the impact of each development activity on code quality. We find that removing the testing activity in the process model results in a significant decrease in Pass@1 accuracy (17.0% to 56.1%). Additionally, eliminating the testing activity leads to a substantial increase in error and warning code smell densities. We also find that the design and code review activities reduce refactor and warning code smells, and improve reliability by adding more exception handling code. Nevertheless, LCG consistently outperforms GPT by reducing code smells and enhancing exception handling. Finally, we find that the GPT model version plays a significant role in the quality of generated code, and LCG helps ensure stability across different versions of LLMs and temperature values.
@@
Originality: We introduce a multi-agent framework called LCG, incorporating software process models from real-world development practice. We integrate agents acting as requirement engineers, architects, developers, testers, and scrum masters, and study how their interaction improves code generation and code quality.
@@
Evaluation: We conduct a fine-grained evaluation on the quality of the generated code using four popular code generation benchmarks: Humaneval [6], Humaneval-ET [8], MBPP [2], MBPP-ET [19], comparing agent interactions and their effect on both accuracy (Pass@1) and other code quality metrics (e.g., smells). We manually checked the generated code and discussed the reasons for test failures. Finally, we examined how different model versions and temperature settings affect code generation stability.
@@
Data Availability: To encourage future research in this area and facilitate replication, We made our data publicly available online https://anonymous.4open.science/r/FlowGen-LLM-E842.
@@ II-A Background
The software development process plays a crucial role in the creation of high-quality software, fundamentally involving communication among various development roles. Given the demonstrated capability of Large Language Model (LLM) agents to mimic domain experts in specific fields [39, 12, 29], this study leverages LLM agents to represent diverse development roles and conduct their associated duties. Our research establishes a collaborative team of LLM agents designed to emulate these process models and roles, aiming to enhance code generation.
@@ II-B Related Works
Code generation is a thriving field of research because of its potential to reduce development costs. Apart from training more powerful LLM to improve code generation, prompt-based and agent-based code generation techniques are two of the most prevalent directions.
Prompt-based Code Generation. Prompt-based code generation employs a range of techniques to refine prompts, ultimately leading to the generation of expected code. For example, Li et al. [18] propose using structured prompts containing code information (e.g., branch and loop structures) to improve the generated code. Nashid et al. [25] retrieval code demos similar to the given task and include them in the prompt to improve code generation. Ruiz et al. [30] use translation techniques for program repair, where buggy programs are first translated into natural language or other programming languages. Then, the translated code is used as a prompt to generate new/fixed code with the same feature. Schäfer et al. [31] iteratively refine prompts based on feedback received from interpreters or test execution results. Kang et al. [15] provide specific instructions, test method signature, and bug report as part of the prompt for generating test code to reproduce bugs. Xie et al. [41] parse the code to identify the focal method and related code context, which are given in the prompt for test code generation. Yuan et al. [48] apply a prompt composition technique by first asking an LLM to provide a high-level description of a method, and then the description is used as part of the prompt to enhance test code generation.
Agent-based Code Generation. Agent-based code generation emphasizes on the importance of role definition and communication among multiple LLM agents. Some approaches incorporate external tools as agents. For example, Huang et al. [14] introduce the test executor agent, employing a Python interpreter to provide test logs for LLMs. Zhong et al. [50] introduces a debugger agent that utilizes a static analysis tool to build control flow graph information, guiding LLMs in locating bugs. Meanwhile, other studies [12, 29, 9] task LLMs as agents by emulating diverse human roles, including analysts, engineers, testers, project managers, chief technology officers (CTOs), etc. These approaches follow the Waterfall model to communicate among these roles, with variation in the prompts and roles, ultimately improving code generation.
@@
While there are many code generation results using the same benchmarks, our evaluation results cannot be directly compared with some other agent-based code generation research [12, 13, 14] due to missing information on the model version, temperature value, steps to post-process the generated code, specific prompts, or how the few-shot samples are selected. Nevertheless, LCG still shows improved Pass@1 compared to existing works that provide clear information. For instance, Dong et al. [9] (uses the same GPT3.5 version as our study) applied few-shot learning with LLM agents and received a Pass@1 of 68.2% in MBPP. In contrast, while uses zero-shot learning, we achieved a Pass@1 of 82.5% in MBPP. We observe the same level of improvements in all the benchmarks. Similarly, Li et al. [18], which used the same model version and temperature value as we do, applied structured chain-of-thought to enhance reasoning ability and achieved a Pass@1 of 60.6% in HumanEval. attained 78.5% under the same model version and temperature setting. In short, due to variations of the LLMs and missing information in prior works, we focus on comparing the LCG’s improvement against GPT (directly using GPT3.5) in this study.
@@ III Methodology
We propose LCG, an agent-based code generation technique based on emulating different software processes. Figure 1 shows the overview of LCG: (1) define the roles and their responsibilities; (2) use LLM agents to represent these roles; and (3) complete the interactions among these agents according to the software process models. In each development activity, we implement chain-of-thought and self-refinement to improve the quality of the generated artifacts. In particular, we study and compare three software process models: Waterfall, TDD, and Scrum. Nevertheless, LCG can be easily adapted to different process models. We use zero-shot learning in all our experiments to avoid biases in selecting data samples. Below, we discuss LCG in detail.
@@ III-A Using LLM Agents to Represent Different Development Roles in Software Process Models
In LCG, we create LLM agents who are responsible for the main development activities: requirement, design, implementation, and testing. Hence, to emulate a software process model, we reorganize the communication and interaction among different agents. The benefit of such a design is that it maximizes the extensibility and reuseable of the agents, and LCG can be easily adapted to different process models.
We implement four agents whose role corresponds to the common development activities: Requirement Engineer, Architect, Developer, and Tester. For Scrum, we introduce an additional role – Scrum Master. We design these roles so that they use the the same prompt template with role-specific details as shown below:
@@
In this prompt template, role refers to one of the roles (e.g., Requirement Engineer) that corresponds to the development activity, and task describes the duties for the role (e.g., analyze and create requirement documents). Instruction leverages chain-of-thought reasoning [38, 18] and refers to role-specific instruction listed in steps, such as 1) analyzing the requirement and 2) writing the requirement documentation. Finally, Context contains the programming question, the agent conversation history, or the agent-generated artifacts. Context includes all necessary information that helps the agents to make a next-step decision based on the current conversation and generated results.
@@
Table I shows the tasks, instructions, and contexts for every development role. In general, every role takes the output from the prior development activities as input (i.e., context). For example, an architect writes a design document based on the requirement document generated by a requirement engineer. We design developers and testers to have multiple tasks. Developers are responsible for writing code and fixing/improving the code based on suggestions. We design testers using a prompt composition technique, which is shown to improve the LLM-generated result [20, 48]. First, testers design a test. Then, testers write and execute the tests based on the design. Finally, testers generate a test failure report. Developers receive the test failure report to fix the code. In addition to the tasks described in Table I, all roles have one common task, which is to provide feedback to other roles for further improvement (e.g., for code review).
Role |
Task |
Instruction |
Context |
Requirement Engineer |
@@
Analyze and generate requirement documentation from the context. |
1) Analyze the requirement; and 2) Write a requirement document. |
Programming problem description. |
Architect |
Design the overall structure and high-level components of the software. |
@@
1) Read the context documents; and 2) Write the design document. The design should be high-level and focus on guiding the developer in writing code. |
Requirement document. |
Developer |
Write code in Python that meets the requirements. |
@@
1) Read the context documents; and 2) Write the code. Ensure that the code you write is efficient, readable, and follows best practices. |
Requirement and design documents. |
Fix the code so that it meets the requirements. |
@@
1) Read the test failure reports and code suggestions from the context; and 2) Rewrite the code. |
Original code, test failure report, and suggestions for improvement. |
Tester |
@@
Design tests to ensure the software satisfies feature needs and quality. |
1) Read context documents; and 2) Design test cases. |
Requirement and design documents. |
Write a Python test script using the unittest framework. |
@@
1) Read the context documents; 2) Write a Python test script; and 3) Follow the input and output given by the requirement. |
Test case design and requirement documents. |
Write a test failure report. |
@@
1) Read the test execution result; and 2) Analyze and generate a test failure report. |
Test execution result. |
Scrum Master |
@@
Summarize and break down the discussion into a task list for the scrum team. |
1) Read and understand the context; and 2) Define the tasks for development roles. |
Meeting discussion. |

### v2

# SOEN-101: Code Generation by Emulating Software Process Models Using Large Language Model Agents

@@ Abstract
Software process models are essential to facilitate collaboration and communication among software teams to solve complex development tasks. Inspired by these software engineering practices, we present FlowGen – a code generation framework that emulates software process models based on multiple Large Language Model (LLM) agents. We emulate three process models, , , and , by assigning LLM agents to embody roles (i.e., requirement engineer, architect, developer, tester, and scrum master) that correspond to everyday development activities and organize their communication patterns. The agents work collaboratively using chain-of-thought and prompt composition with continuous self-refinement to improve the code quality. We use GPT3.5 as our underlying LLM and several baselines (RawGPT, CodeT, Reflexion) to evaluate code generation on four benchmarks: HumanEval, HumanEval-ET, MBPP, and MBPP-ET. Our findings show that excels compared to other process models, achieving a Pass@1 of 75.2, 65.5, 82.5, and 56.7 in HumanEval, HumanEval-ET, MBPP, and MBPP-ET, respectively (an average of 15% improvement over RawGPT). Compared with other state-of-the-art techniques, achieves a higher Pass@1 in MBPP compared to CodeT, with both outperforming Reflexion. Notably, integrating CodeT into resulted in statistically significant improvements, achieving the highest Pass@1 scores. Our analysis also reveals that the development activities impacted code smell and exception handling differently, with design and code review adding more exception handling and reducing code smells. Finally, FlowGen models maintain stable Pass@1 scores across GPT3.5 versions and temperature values, highlighting the effectiveness of software process models in enhancing the quality and stability of LLM-generated code.
@@
Given the complexity of software development, LLM agents stand out among various LLM techniques. Agents are LLM instances that can be customized to carry out specific tasks that replicate human workflow [14, 11]. Recently, multi-agent systems have achieved significant progress in solving complex problems in software development by emulating development roles [14, 11, 31]. MetaGPT, introduced by Hong et al. [14], integrated development workflow using standard operating procedures by assigning specific roles (e.g., a designer or a developer) to LLM agents. Dong et al. [11] developed self-collaboration, which assigns LLM agents to work as distinct “experts” for sub-tasks in software development. Qian et al. [31] proposed an end-to-end framework for software development through self-communication among the agents.
@@
Despite the promising applications of LLMs in automating software engineering tasks, it is pivotal to recognize that software development is a collaborative and multi-faceted endeavor. In practice, developers and stakeholders work together, following certain software process models like Waterfall, Test-Driven-Development (TDD), and Scrum. The process models help facilitate communication and collaboration to ensure the delivery of high-quality products. Even though there is a common community agreement on the pros and cons of each process model [13], the impact of adopting these process models for LLM code generation tasks remains unknown. In particular, will emulating different process models impact the generated code quality in different aspects, such as reliability, code smell, and functional correctness?
@@
While some research has explored integrating multi-agents within LLM frameworks [31, 45, 47], their research focus diverges from the influence of the software process model on code generations for several reasons: 1) Xu et al. [45] do not adhere to specific process models, and 2) both Dong et al. [11] and Qian et al. [31] focus solely on Waterfall-like models, neglecting TDD and Scrum, which may have different impact on code generations. Importantly, none of the aforementioned studies conduct a fine-grain analysis of how different development activities affect code quality metrics, such as code smell and reliability, other than the Pass@1 score. Our study takes further steps to analyze the impacts of different agents within the process models on code generation and their influence on other code quality attributes.
@@
This paper presents a novel multi-agent LLM-based code generation framework named FlowGen. FlowGen integrates diverse prompt engineering techniques, including chain-of-thought [40, 20], prompt composition [22, 50], and self-refinement [25], with a focus on emulating the flow of development activities in various software process models. Specifically, we implemented three popular process models into FlowGen: , , and . Each process model emulates a real-world development team involving several LLM agents whose roles (i.e., requirement engineer, architect, developer, tester, and scrum master) correspond to common software development activities. The agents work collaboratively to produce software artifacts and help other agents review and improve the artifacts in every activity.
@@
We evaluate FlowGen on four popular code generation benchmarks: HumanEval [8], HumanEval-ET [10], MBPP [3], and MBPP-ET [21]. We apply zero-shot learning to avoid biases in selecting few-shot samples [44]. To compare, we also apply zero-shot learning on GPT-3.5 as our baseline (RawGPT). We repeat our experiments five times to account for variability in LLM’s responses and report the average value and standard deviation. To study code quality, in addition to Pass@1, we run static code checkers to detect the prevalence of code smells in the generated code. Our evaluation shows that ’s generated code achieves the highest accuracy (Pass@1 is 75.2 for HumanEval, 65.5 for HumanEval-ET, 82.5 for MBPP, and 56.7 for MBPP-ET), surpassing RawGPT’s Pass@1 by 5.2% to 31.5%. While FlowGen, in general, is more stable than RawGPT, exhibits the most stable results with an average standard deviation of only 1.3% across all benchmarks.
Additionally, we compare with other state-of-the-art techniques: CodeT [5] and Reflexion [36]. Both and CodeT outperform Reflexion significantly in terms of Pass@1 across all benchmarks, with achieving a higher Pass@1 than CodeT in MBPP. Furthermore, the integration of CodeT into demonstrates the highest Pass@1 scores, highlighting the potential of integrating other prompting techniques with FlowGen for improved code generation.
@@
We further study the impact of each development activity on code quality. We find that removing the testing activity in the process model results in a significant decrease in Pass@1 accuracy (17.0% to 56.1%). Eliminating the testing activity also leads to a substantial increase in error and warning code smell densities. We also find that the design and code review activities reduce refactor and warning code smells, and improve reliability by adding more exception handling code. Nevertheless, FlowGen consistently outperforms RawGPT by reducing code smells and enhancing exception handling. Finally, we find that the GPT model version plays a significant role in the quality of generated code, and FlowGen helps ensure stability across different versions of LLMs and temperature values.
@@
Originality: We introduce a multi-agent framework called FlowGen, incorporating software process models from real-world development practice. We integrate agents acting as requirement engineers, architects, developers, testers, and scrum masters, and study how their interaction improves code generation and code quality.
@@
Evaluation: We conduct a fine-grained evaluation on the quality of the generated code using four popular code generation benchmarks: Humaneval [8], Humaneval-ET [10], MBPP [3], MBPP-ET [21], comparing agent interactions and their effect on both accuracy (Pass@1) and other code quality metrics (e.g., smells). We manually checked the generated code and discussed the reasons for test failures. Finally, we examined how model versions and temperature settings affect code generation stability.
@@
Data Availability: To encourage future research in this area and facilitate replication, We made our data and code publicly available online [2].
@@ II-A Background
The software development process plays a crucial role in software development, fundamentally involving communication among various development roles. Given the demonstrated capability of Large Language Model (LLM) agents to mimic domain experts in specific fields [41, 14, 31], this study leverages LLM agents to represent diverse development roles and conduct their associated duties. Our research establishes a collaborative team of LLM agents designed to emulate these process models and roles, aiming to enhance code generation.
@@ II-B Related Works
Code generation is a thriving field of research because of its potential to reduce development costs. In particular, prompt-based and agent-based code generation techniques are two of the most prevalent directions.
Prompt-based Code Generation. Prompt-based code generation employs a range of techniques to refine prompts, ultimately leading to the generation of expected code. For example, Li et al. [20] propose using structured prompts containing code information (e.g., branch and loop structures) to improve the generated code. Nashid et al. [27] retrieval code demos similar to the given task and include them in the prompt to improve code generation. Ruiz et al. [32] use translation techniques for program repair, where buggy programs are first translated into natural language or other programming languages. The translated code is used as a prompt to generate new/fixed code with the same feature. Schäfer et al. [33] iteratively refine prompts based on feedback received from interpreters or test execution results. Kang et al. [17] provide specific instructions, test method signature, and bug report as part of the prompt for generating test code to reproduce bugs. Xie et al. [43] parse the code to identify the focal method and related code context, which are given in the prompt for test code generation. Yuan et al. [50] apply a prompt composition technique by first asking an LLM to provide a high-level description of a method, and then the description is used as part of the prompt to enhance test code generation. Chen et al. [5] introduced CodeT, a framework that employs self-generated tests to evaluate the quality of generated code. Shinn et al. [36] presented Reflexion, which utilizes an evaluator LLM to provide feedback for enhancing future decision-making processes.
Agent-based Code Generation. Agent-based code generation emphasizes on the importance of role definition and communication among multiple LLM agents. Some approaches incorporate external tools as agents. For example, Huang et al. [16] introduce the test executor agent, employing a Python interpreter to provide test logs for LLMs. Zhong et al. [52] introduces a debugger agent that utilizes a static analysis tool to build control flow graph information, guiding LLMs in locating bugs. Meanwhile, other studies [14, 31, 11] task LLMs as agents by emulating diverse human roles, including analysts, engineers, testers, project managers, chief technology officers (CTOs), etc. Nevertheless, these studies miss key roles in the development activities (e.g., only has analysts, coders, and testers [11]) or focus more on the business side of the roles (e.g., employ CTO and CEO) [31]. In our work, we try to follow the Waterfall model that is proposed in the software engineering literature and create agents that correspond to every development activity. These approaches follow the Waterfall model to communicate among these roles, with variation in the prompts and roles, ultimately improving code generation.
@@ III Methodology
We propose FlowGen, an agent-based code generation technique based on emulating different software processes. Figure 1 shows the overview of FlowGen: (1) define the roles and their responsibilities; (2) use LLM agents to represent these roles; and (3) complete the interactions among these agents according to the software process models. In each development activity, we implement chain-of-thought and self-refinement to improve the quality of the generated artifacts. In particular, we study and compare three software process models: Waterfall, TDD, and Scrum. Nevertheless, FlowGen can be easily adapted to different process models. We use zero-shot learning in all our experiments to avoid biases in selecting data samples. Below, we discuss FlowGen in detail.
@@ III-A Using LLM Agents to Represent Different Development Roles in Software Process Models
In FlowGen, we create LLM agents who are responsible for the main development activities: requirement, design, implementation, and testing. Hence, to emulate a software process model, we reorganize the communication and interaction among different agents. The benefit of such a design is that it maximizes the extensibility and reusability of the agents, and FlowGen can be easily adapted to different process models. We implement four agents whose role corresponds to the common development activities: Requirement Engineer, Architect, Developer, and Tester. For Scrum, we introduce an additional role – Scrum Master.
We designed these roles to use the same prompt template across different process models (with different terms such as user stories v.s. requirement) to investigate the effectiveness of process models on code generation. The exact words that we used for the prompts can be found online [2]. The role-specific details of our prompt are:
@@
In this prompt template (inspired by MetaGPT [14] and Self-Collaboration [11]), role refers to one of the roles (e.g., Requirement Engineer) that corresponds to the development activity, and task describes the duties for the role (e.g., analyze and create requirement documents). Instruction leverages chain-of-thought reasoning [40, 20] and refers to role-specific instruction listed in steps, such as 1) analyzing the requirement and 2) writing the requirement documentation. Finally, Context contains the programming question, the agent conversation history, or the agent-generated artifacts. Context includes all necessary information that helps the agents to make a next-step decision based on the current conversation and generated results.
@@
Table I shows the tasks, instructions, and contexts for every development role. In general, every role takes the output from the prior development activities as input (i.e., context). For example, an architect writes a design document based on the requirement document generated by a requirement engineer. We design developers and testers to have multiple tasks. Developers are responsible for writing code and fixing/improving the code based on suggestions. We design testers using a prompt composition technique, which is shown to improve the LLM-generated result [22, 50]. First, testers design a test. Then, testers write and execute the tests based on the design. On average, FlowGen generates four tests for each problem before the review meeting and six after the meeting. It is important to note that oracles are kept aside and never used in the code-generation process. Finally, testers generate a test failure report.
Developers receive the test failure report to fix the code. In addition to the tasks described in Table I, all roles have one common task, which is to provide feedback to other roles for further improvement (e.g., for code review).
| Role | Task | Instruction | Context |
@@
| Requirement Engineer | Analyze and generate requirement documentation from the context. | 1) Analyze the requirement; and 2) Write a requirement document. | Programming problem description. |
@@
| Architect | Design the overall structure and high-level components of the software. | 1) Read the context documents; and 2) Write the design document. The design should be high-level and focus on guiding the developer in writing code. | Requirement document. |
@@
| Developer | Write code in Python that meets the requirements. | 1) Read the context documents; and 2) Write the code. Ensure that the code you write is efficient, readable, and follows best practices. | Requirement and design documents. |
@@
| Fix the code so that it meets the requirements. | 1) Read the test failure reports and code suggestions from the context; and 2) Rewrite the code. | Original code, test failure report, and suggestions for improvement. | |
@@
| Tester | Design tests to ensure the software satisfies feature needs and quality. | 1) Read context documents; and 2) Design test cases. | Requirement and design documents. |
@@
| Write a Python test script using the unittest framework. | 1) Read the context documents; 2) Write a Python test script; and 3) Follow the input and output given by the requirement. | Test case design and requirement documents. | |
@@
| Write a test failure report. | 1) Read the test execution result; and 2) Analyze and generate a test failure report. | Test execution result. | |
@@
| Scrum Master | Summarize and break down the discussion into a task list for the scrum team. | 1) Read and understand the context; and 2) Define the tasks for development roles. | Meeting discussion. |